OpenStudy (anonymous):

Help me with this please . Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set # 1 --> (1, 5) and (2, 9) Set # 2 --> (6, 2) and (3, 4

OpenStudy (anonymous):

The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (anonymous):

First thing you need to do is find the slope of each set. Use the equation y-y/x-x. Make sure you subtract the first y from the second y and the first x from the second x. Set #1 9-5/2-1=4 Set#2 4-2/3-6=-2/3 With a slope of 4 and then a slope of -2/3, the second statement is true.